Anna flies back and forth between 2 cities that are 1680 miles apart. It takes 4 hours to fly going into the wind and 3.5 hours with the wind. Find the wind speed and the speed of the plane in still air.
Thanks!

Answer by checkley77(12844)

1,680/4=420 mph.
1,680/3.5=480 mph.
(480-420)/2=60/2=30 mph. is the speed of the wind.
480-30=450 mph. is the speed of the plane in still air.
420+30=450 ditto.